Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labor-hours and its standard cost card per unit is as follows:

Direct materials: 4 pounds at $8 per pound $ 32
Direct labor: 2 hours at $16 per hour 32
Variable overhead: 2 hours at $6 per hour 12
Total standard cost per unit $ 76

The planning budget for March was based on producing and selling 32,000 units. However, during March the company actually produced and sold 37,000 units and incurred the following costs:

  1. Purchased 160,000 pounds of raw materials at a cost of $7.40 per pound. All of this material was used in production.
  2. Direct laborers worked 67,000 hours at a rate of $17 per hour.

  3. The total variable manufacturing overhead for the month was $422,100.

QUESTION:

What is the labor efficiency variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input all amounts as positive values.)
Answer #1

Labor efficiency variance

= (SH-AH) * SR

= (37,000*2 - 67,000) * 16

= (74,000 - 67,000) * 16

= 112,000 Favourable